OpenStudy (anonymous):

ALGEBRAIC EXPRESSION the product of 5 less than a number and 10

OpenStudy (anonymous):

product = multiply less than = subtract, but you switch around the the order. (example: 8 less than x means x - 8) "a number" = means that you need to give any variable you want. a, b, c...x is the most common. answer: (x-5) * 10

OpenStudy (anonymous):

Let x = 'a number' x - 5 = '5 less than a number' (x-5) * 10 = product of 5 less than a number and 10 So that's your Algebraic Expression

OpenStudy (anonymous):

I got the same thing unless you left out a part of a question making it an equation which in that case is solvable

OpenStudy (anonymous):

Just looking for an expression, not necessarily an equation